OpenStudy (anonymous):

Farmer Joe wants to fence an area of 1.4 million square feet in a rectangular field and then divide it in half with a fence parallel to one of the sides of the rectangular. What is the least amount of fencing required if the partitian is : a) Horizontal b) Vertical

OpenStudy (anonymous):

Let w be the width of the field, and l be the length w * l = 1, 400, 000 So he has 2w + 2l feet of fence to fence that field Then, if he halves that field with a horizontal fence, he uses another l feet of fence. So he uses 2w + 3l feet of fence. Let F be the fencing required: F = 2w + 3l But l = 1,400,000/w, so: F = 2w + 4,200,000/w dF/dw = 2 - 4,200,000/w^2 We find the minimum value of F when dF/dw = 0: 2 - 4,200,000/w^2 = 0 Hence 2w^2 = 4,200,000 Hence w^2 = 2,100,000 Hence w = 1.45 * 10^3 gives the minimum, and the value of F at that minimum is 5.8 * 10^3 feet Thus for a horizontal partitian, the least amount of fencing is roughly 5,800 feet. For a vertical partitian: F = 3w + 2l Hence F = 3w + 2,800,000/w So dF/dw = 3 - 2,800,000/w^2 This is equal to 0 when 3w^2 = 2,800,000 so when w = 966.0918 The value of F when w = 966.0918 is given by 5.8 * 10^3 feet So the least amount of fencing required is the same for both vertical and horizontal partitians, so it doesn't matter how the farmer should divide the field in half.